# On the Duality between Gradient Transformations and Adapters

### v1

@@ 4 Empirical Study
The equivalences in § 3 are agnostic to the choice of left and right transformations in ${\mathbf{S}}={\mathbf{R}}^{\top}\otimes{\mathbf{L}}$. However, one might expect that the choice of ${\mathbf{L}}$ and ${\mathbf{R}}$ should matter for downstream performance. Hence, in the following sections, we first explore how the choice of ${\mathbf{S}}$ affects pretraining [^5] We target the pretraining setting as the gap between ordinary training and memory-efficient training methods is typically larger in pretraining than it is in finetuning. performance, and how by viewing gradient transformations as adapters, we further improve memory efficiency by combining the technique with QLoRA-style (Dettmers et al., 2023) training (§ 4.1). We then show how the converse is also useful: by viewing LoRA adapters through the lens of gradient transformations, we can improve distributed training of LoRA adapters by coordinating the LoRA adapter initialization across different workers (§ 4.2).
Experimental setup. We consider two moderate-scale language modeling settings: a 200M setting (training on 5B tokens) and a 1.3B setting (training on 10B tokens). [^6] While this is not large by modern standards, due to our limited compute resources this is the largest setting at which we can feasibly perform experiments. We use a Transformer++ (Touvron et al., 2023a) architecture and train on the SlimPajama (Soboleva et al., 2023) dataset, tokenized using the Llama-2 (Touvron et al., 2023b) tokenizer, using sequences of length 2048. All numbers we report are perplexity on a disjoint (validation) set of SlimPajama. We use AdamW (Loshchilov & Hutter, 2019) with weight decay $0.1$, $\beta_{1}=0.9$ and $\beta_{2}=0.95$. We warm up the learning rate to $4\times 10^{-4}$, before decaying it via a cosine decay schedule to $1\times 10^{-4}$. We conduct all training in bfloat16 precision. See App. C for more details.
@@ 4.1 Study 1: Memory-Efficient Pretraining
The discussion in § 3 establishes a direct link between GaLore and one-sided LoRA. But how should we set ${\mathbf{S}}$ in practice? From the perspective of accurate gradient estimation, it would be ideal to have ${\mathbf{S}}^{\top}{\mathbf{S}}\approx{\mathbf{I}}$, since in the vanilla SGD case this would be equivalent to performing SGD with sketched gradients, where ${\mathbf{S}}^{\top}{\mathbf{S}}\overline{\Theta}\approx\overline{\Theta}$ (Murray et al., 2023). For the GaLore case with ${\mathbf{S}}={\mathbf{I}}\otimes{\mathbf{P}}$, this amounts to setting ${\mathbf{P}}$ such that ${\mathbf{P}}{\mathbf{P}}^{\top}\approx{\mathbf{I}}$, which could be achieved by, e.g., using random sketching matrices with the property $\mathbb{E}[{\mathbf{P}}{\mathbf{P}}^{\top}]=\mathbf{I}$ [^7] This sketching view of LoRA provides a possible perspective on why one-sided LoRA finetuning works well in practice (Zhang et al., 2023; Zhu et al., 2024; Hayou et al., 2024). As noted by Hao et al. (2024), using a random sketching matrix can enable further savings as only the seed needs to be persisted across optimization steps. We thus experiment with a variety of sketching matrices for LoRA-based pretraining as shown in Tab. 2.
Another benefit of the adapter parameterization of gradient projections is that it allows us to be more memory efficient by quantizing the base weights as done in QLoRA (Dettmers et al., 2023). Specifically, given the adapter parameterization $\Theta+{\mathbf{S}}^{\top}\Lambda$ we can quantize $\Theta$ and only train $\Lambda$, thus enabling further memory savings. Finally, the adapter parameterization has the additional benefit of reducing the number of trainable parameters being registered for automatic differentiation, which allows for gradient accumulation to happen in a lower dimensional space. (See App. C for more discussion.)
@@
Results. The results are shown in Tab. 2, where we follow the original GaLore paper and use a rank of $256$ for the 200M model and a rank of $512$ for the 1.3B model, [^8] The only exceptions are for the double-sided methods. For the two-sided Gaussian, we set the rank as to match the number of trainable parameters in the one-sided Gaussian approach. For the two-sided SVD, we use the same rank as in two-sided Gaussian, which incurs more memory since the projection matrices must be persisted across optimization steps. and further merge the adapters into the full weights and reinitialize them every 200 steps. We see that one-sided transformations, regardless of their nature, perform somewhat similarly at both 200M and 1.3B scale, suggesting that using a random gradient transformation matrix may be a more economical alternative to using the top singular vectors derived from the gradient as in GaLore. We also find that ReLoRA performs comparably to one-sided gradient transformations, suggesting that the additional flexibility of ReLoRA (i.e., optimizing two sides of a LoRA adapter instead of only one side) is not necessary. Interestingly, using two-sided Gaussian gradient transformations, which are similar in spirit to recent approaches for memory-efficient finetuning (Jiang et al., 2024; Bałazy et al., 2024; Wang et al., 2025), degrades performance when memory consumption is matched to one-sided methods; two sided SVD-based projections fare slightly better but still trail behind one-sided methods and incur a much larger memory cost, since two projection matrices must be persisted. While Zhao et al. (2024) report no gap between GaLore and full pretraining, we did not find this to be true on our setup, [^9] Which is different from theirs in many ways, e.g., we train on longer sequences with gradient accumulation using bfloat16. and instead observe a non-trivial gap between regular (full) training and these memory-efficient pretraining methods.
@@
Figure 1: Average gradient reconstruction error $\frac{1}{d}\|\overline{\Theta}-{\mathbf{S}}^{\top}{\mathbf{S}}\overline{\Theta}\|_{2}^{2}$ of the various transformations across training steps at 200M scale.
@@
When adding quantization to the base weights (where we use groups of size 256), we find that, across the board, 8-bit integer quantization can be performed without major performance degradation, whereas 4-bit NormalFloat quantization begins to incur a penalty (4-bit integer did even worse). Finally, we find that QGaLore (Zhang et al., 2024), which quantizes the weights to INT8 precision and trains these INT8 weights directly using an SVD gradient transformation, underperforms QLoRA-style approach to quantized GaLore training.
@@
Analysis. We have motivated our experiments with sketching matrices from the perspective of accurate gradient compression, i.e., we use ${\mathbf{S}}$ to compress the gradient, and then ${\mathbf{S}}^{\top}$ to decompress it. From this compression viewpoint, one may then wonder whether different gradient transformations exhibit different reconstruction capabilities, and whether this ultimately dictates the performance of the resulting model. As shown in Fig. 1, we find that the gradient reconstruction error does not correlate with performance. As expected, methods that perform SVD on the gradients have low reconstruction error (since SVD explicitly minimizes a reconstruction objective), but as shown in Tab. 2, SVD performs similarly to sketching matrices, which have higher reconstruction error. We believe that the relationship between the nature of the gradient transformation and downstream performance is fairly complex, and merits further investigation. Fig. 2 (appendix) shows similar results for cosine similarity instead of squared error.
@@ 4.2 Study 2: Distributed Pretraining
Dist. Training (DiLoCo) -- 18.000 189 5818.0001895818.000\,189\,5818.000 189 58 12.767 859 412.767859412.767\,859\,412.767 859 4
Dist. ReLoRA (LTE) -- 20.973 187 5520.9731875520.973\,187\,5520.973 187 55 13.715 583 3613.7155833613.715\,583\,3613.715 583 36
Identical Random ${\mathbf{P}}_{i}={\mathbf{P}}_{j}$ 21.507 394 5121.5073945121.507\,394\,5121.507 394 51 14.279 870 7914.2798707914.279\,870\,7914.279 870 79
Independent Random $\operatorname{\mathbb{E}}[{\mathbf{P}}_{i}{\mathbf{P}}_{j}^{\top}]=\mathbf{0}$ 20.114 808 0120.1148080120.114\,808\,0120.114 808 01 13.655 087 1713.6550871713.655\,087\,1713.655 087 17
Distributed Random ${\mathbf{P}}_{i}{\mathbf{P}}_{j}^{\top}=\mathbf{0}$ 19.813 587 5219.8135875219.813\,587\,5219.813 587 52 13.514 498 9213.5144989213.514\,498\,9213.514 498 92
@@
Our second experiment targets distributed pretraining of LLMs across poorly-connected and resource-constrained workers, which is important for many applications of interest, from federated training of LLMs to scaling up LLMs across data centers that are not co-located, i.e., where techniques like FSDP are not possible. DiLoCo (Douillard et al., 2024) is a recent effective approach that has workers train independently for some number of iterations using an inner optimizer, then uses the average change in parameters from each worker as a “pseudo-gradient” on an outer optimizer that updates a global copy of the parameters (i.e., as in federated learning; McMahan et al., 2017; Reddi et al., 2021). This approach has since been scaled up to train 10B LLMs across distributed workers (Jaghouar et al., 2024).
However, DiLoCo still assumes that each worker has enough memory to perform a full forward/backward pass on the model, i.e., it does not target memory efficiency. A memory-efficient distributed approach that is of particular interest in light of the equivalence in § 3 is LoRA-the-explorer (LTE; Huh et al., 2024), which can be seen as an extension of ReLoRA to the distributed setting. LTE has $K$ independent workers train separate LoRA adapters for a small number of local steps, and then performs a global step by averaging the adapters across workers. The globally-averaged adapter is then merged into the base weights, and optimization continues by resetting and traiing the worker-specific LoRA adapters.
We will now describe how the equivalence in § 3 can be used to derive an improved version of LTE, which trains only one side of the LoRA adapter in each worker, but initializes the frozen side in a worker-aware manner. Consider a one-sided analogue of LTE, where the weight ${\mathbf{W}}^{(g,l)}_{k}$ for the $k$th worker after $g$ global and $l$ local updates is
@@
By § 3.2, local steps must correspond to training the worker weights using a gradient transformation,
@@
One approach to initialize/reset the frozen side of worker-specific LoRA adapters (i.e., the gradient projections ${\mathbf{P}}_{k}$) is to sample a projection and broadcast it to all workers. However, the GaLore–LoRA duality suggests a different scheme. § 3.1 shows that training with linear gradient transformations only optimizes a subspace of the full model, namely $\operatorname{range}({\mathbf{S}}^{\top})$ where ${\mathbf{S}}^{\top}={\mathbf{I}}\otimes{\mathbf{P}}^{\top}$ in the GaLore case. This suggests a different approach to distributed LoRA training, wherein the frozen part of each LoRA adapter, ${\mathbf{S}}_{1},\ldots,{\mathbf{S}}_{K}$, is initialized differently, so that the sum of their ranges allows a larger subspace to be trained. For example, we could sample random semi-orthogonal matrices ${\mathbf{P}}_{1},\ldots,{\mathbf{P}}_{K}$ uniformly at random, assign each worker ${\mathbf{S}}_{i}={\mathbf{I}}\otimes{\mathbf{P}}^{\top}_{i}$, and they will likely each cover different portions of the space. An even stronger strategy would be to demand that $\operatorname{range}({\mathbf{S}}^{\top}_{1}),\ldots,\operatorname{range}({\mathbf{S}}^{\top}_{K})$ must be orthogonal, which can be realized by keeping the ${\mathbf{P}}_{i}$’s as semi-orthogonal, but enforcing that ${\mathbf{P}}_{i}^{\top}{\mathbf{P}}_{j}={\mathbf{I}}$ (e.g., by generating a random $m\times m$ orthogonal matrix and having each worker take a different $d\times m$ submatrix.) Intuitively, this ensures that no worker is duplicating the work of another, since their projections are pairwise orthogonal. We experiment with such identical random, independent random, distributed random initialization schemes.
@@
Dist. Training (DiLoCo) 17.809 353 3317.8093533317.809\,353\,3317.809 353 33 18.000 189 5818.0001895818.000\,189\,5818.000 189 58 18.563 059 7818.5630597818.563\,059\,7818.563 059 78
Dist. ReLoRA (LTE) 23.761 633 0623.7616330623.761\,633\,0623.761 633 06 20.973 187 5520.9731875520.973\,187\,5520.973 187 55 19.538 457 2719.5384572719.538\,457\,2719.538 457 27
Identical Random 23.955 952 7223.9559527223.955\,952\,7223.955 952 72 21.507 394 5121.5073945121.507\,394\,5121.507 394 51 20.316 635 6120.3166356120.316\,635\,6120.316 635 61
Independent Random 20.643 903 4620.6439034620.643\,903\,4620.643 903 46 20.114 808 0120.1148080120.114\,808\,0120.114 808 01 19.971 539 4619.9715394619.971\,539\,4619.971 539 46
Distributed Random 20.317 105 4720.3171054720.317\,105\,4720.317 105 47 19.813 587 5219.8135875219.813\,587\,5219.813 587 52 19.655 980 3219.6559803219.655\,980\,3219.655 980 32
@@
Table 4: Results of the distributed pretraining experiments as we vary the rank of the gradient projections and number of workers. For the DiLoCo baseline, we only vary the number of workers; note that the for the distributed ReLoRA baseline (which is similar to LTE, Huh et al., 2024), we have double the number of trainable parameters as in the one-sided methods.
@@
Results. The results for the main set of experiments are shown in Tab. 3. We consider two baselines: (i) DiLoCo (Douillard et al., 2024), which has each worker training independently for $500$ steps before computing a pseudo-gradient that is used to update the global parameters using SGD with Nesterov momentum (Sutskever et al., 2013), and (ii) distributed ReLoRA, which is an analog of ReLoRA but adapted to train like DiLoCo, i.e., one trains the LoRA adapter for $500$ steps and defines the adapter weight as the pseudo-gradient for the Nesterov step; this is very close to LTE. [^10] LTE can be seen as using SGD as the optimizer on the pseudo-gradients, but we found this led to worse results in preliminary experiments. Our distributed GaLore experiments make use of random semi-orthogonal projections since the distributed random initialization for it is easy to compute, and does not add significant communication overhead. [^11] Each worker just needs the seed used to sample the orthogonal matrix, and the indices of the rows it will keep. As in § 4.1, distributed GaLore leads to degradations at both 200M and 1.3B scales compared to the full distributed training baseline (i.e., DiLoCo). However, our distributed random initialization scheme, where workers are “aware” of each, performs well, thus demonstrating the utility of the gradient transformation–adapter duality from § 3.
@@
Analysis. We perform a study at the 200M scale over how the number of workers and rank affect performance. Intuitively, larger ranks lead to a larger subspace being trained by each worker (and, in the limit, we should recover something akin to DiLoCo when there is no rank reduction), so we would expect performance to improve as we increase the rank. Indeed, the results for this ablation (shown in Tab. 4) confirm this intuition, likely because DiLoCo benefits from more workers to get a better estimate of the pseudo-gradient for the outer optimizer step. More surprisingly, we find that this gap is largely bridged by ensuring that different gradient transformations are assigned to each worker, with the distributed initialization once again performing the best. It would be interesting to further study how the effectiveness of the distributed initialization scheme changes as we go to more extreme settings (e.g., hundreds of extremely low-rank workers).

### v2

@@ 4 Empirical Study
The equivalences in §˜3 are agnostic to the choice of left and right transformations in ${\mathbf{S}}={\mathbf{R}}^{\top}\otimes{\mathbf{L}}$. However, one might expect that the choice of $\mathbf{L}$} and $\mathbf{R}$} should matter for downstream performance. Hence, in the following sections, we first explore how the choice of $\mathbf{S}$} affects pretraining [^6] We target the pretraining setting as the gap between ordinary training and memory-efficient training methods is typically larger in pretraining than it is in finetuning. performance, and how by viewing gradient transformations as adapters, we further improve memory efficiency by combining the technique with QLoRA-style (Dettmers et al., 2023) training (§˜4.1). We then show how the converse is also useful: by viewing LoRA adapters through the lens of gradient transformations, we can improve distributed training of LoRA adapters by coordinating the LoRA adapter initialization across different workers (§˜4.2).
Experimental setup. We consider two moderate-scale language modeling settings: a 200M setting (training on 5B tokens) and a 1.3B setting (training on 10B tokens). [^7] While this is not large by modern standards, due to our limited compute resources this is the largest setting at which we can feasibly perform experiments. We use the Llama Transformer architecture (Touvron et al., 2023a) and train on the SlimPajama (Soboleva et al., 2023) dataset, tokenized using the Llama-2 (Touvron et al., 2023b) tokenizer, using sequences of length 2048. All numbers we report are perplexity on a disjoint (validation) set of SlimPajama. We use AdamW (Loshchilov & Hutter, 2019) with weight decay $0.1$, $\beta_{1}=0.9$ and $\beta_{2}=0.95$. We warm up the learning rate to $4\times 10^{-4}$, before decaying it via a cosine decay schedule to $1\times 10^{-4}$. We conduct all training in bfloat16 precision. See App.˜D for more details on our experimental setup.
@@ 4.1 Study 1: Memory-Efficient Pretraining
The discussion in §˜3 establishes a direct link between GaLore and one-sided LoRA. But how should we set $\mathbf{S}$} in practice? From the perspective of accurate gradient estimation, it would be ideal to have ${\mathbf{S}}^{\top}{\mathbf{S}}\approx{\mathbf{I}}$, since in the vanilla SGD case this would be equivalent to performing SGD with sketched gradients, where ${\mathbf{S}}^{\top}{\mathbf{S}}\overline{\Theta}\approx\overline{\Theta}$ (Murray et al., 2023). For the GaLore case with ${\mathbf{S}}={\mathbf{I}}\otimes{\mathbf{P}}$, this amounts to setting $\mathbf{P}$} such that ${\mathbf{P}}{\mathbf{P}}^{\top}\approx{\mathbf{I}}$, which could be achieved by, e.g., using random sketching matrices with the property $\mathbb{E}[{\mathbf{P}}{\mathbf{P}}^{\top}]=\mathbf{I}$. [^8] This sketching view of LoRA provides a possible perspective on why one-sided LoRA finetuning works well in practice (Zhang et al., 2023; Zhu et al., 2024; Hayou et al., 2024). As noted by Hao et al. (2024), using a random sketching matrix can enable further savings as only the random number generator (RNG) seed needs to be persisted across optimization steps. We thus experiment with a variety of sketching matrices for LoRA-based pretraining as shown in Tab.˜1.
Another benefit of the adapter parameterization of gradient projections is that it allows us to be more memory efficient by quantizing the base weights as done in QLoRA (Dettmers et al., 2023). Specifically, given the adapter parameterization $\Theta+{\mathbf{S}}^{\top}\Lambda$ we can quantize $\Theta$ and only train $\Lambda$, thus enabling further memory savings. Finally, the adapter parameterization has the additional benefit of reducing the number of trainable parameters being registered for automatic differentiation, which allows for gradient accumulation to happen in a lower dimensional space. (See App.˜D for more discussion.)
@@
Results. The results are shown in Tab.˜2, where we follow the original GaLore paper and use a rank of $256$ for the 200M model and a rank of $512$ for the 1.3B model, [^9] The only exceptions are for the double-sided methods. For the two-sided Gaussian, we set the rank as to match the number of trainable parameters in the one-sided Gaussian approach. For the two-sided SVD, we use the same rank as in two-sided Gaussian, which incurs more memory since the projection matrices must be persisted across optimization steps. and further merge the adapters into the full weights and reinitialize them every 200 steps. We see that one-sided transformations, regardless of their nature, perform somewhat similarly at both 200M and 1.3B scale, suggesting that using a random gradient transformation matrix that can be cheaply rematerialized on-the-fly may be more economical than using the top singular vectors derived from the gradient as in GaLore. We also find that ReLoRA performs comparably to one-sided gradient transformations, suggesting that the additional flexibility of ReLoRA (i.e., optimizing two sides of a LoRA adapter instead of only one side) is not necessary. Using two-sided Gaussian gradient transformations degrades performance when memory consumption is matched to one-sided methods; two-sided SVD-based projections fare slightly better but still trail behind one-sided methods and incur a much larger memory cost, since two projection matrices must be persisted. While Zhao et al. (2024) report no gap between GaLore and full pretraining, we did not find this to be true on our setup, [^10] Which is different from theirs in many ways, e.g., we train on longer sequences with gradient accumulation using bfloat16. and instead observe a non-trivial gap between regular (full) training and these memory-efficient pretraining methods.
@@
Figure 1: Average of gradient reconstruction error $\|\overline{\Theta}-{\mathbf{S}}^{\top}{\mathbf{S}}\overline{\Theta}\|_{2}^{2}$ of the various transformations across training steps at 200M scale.
@@
When adding quantization to the base weights (where we use groups of size 256), we find that, across the board, 8-bit integer quantization can be performed without major performance degradation, whereas 4-bit NormalFloat quantization begins to incur a penalty (4-bit integer did even worse). This suggests that adapter-based training, with a rematerializable gradient transformation and 8-bit integer quantization of the base weights, is a promising recipe for memory-efficient pretraining. Finally, we find that QGaLore (Zhang et al., 2024), which quantizes the weights to INT8 precision and trains these INT8 weights directly using an SVD gradient transformation, underperforms a QLoRA-style approach to quantized GaLore training.
@@
Analysis. We have motivated our experiments with sketching matrices from the perspective of accurate gradient compression, i.e., we use $\mathbf{S}$} to compress the gradient, and then $\mathbf{S}$$^{\top}$ to decompress it. From this compression viewpoint, one may then wonder whether different gradient transformations exhibit different reconstruction capabilities, and whether this ultimately dictates the performance of the resulting model. As shown in Fig.˜1, we find that the gradient reconstruction error does not correlate with performance. As expected, methods that perform SVD on the gradients have low reconstruction error (since SVD explicitly minimizes a reconstruction objective), but as shown in Tab.˜2, SVD performs similarly to sketching matrices, which have higher reconstruction error. We believe that the relationship between the nature of the gradient transformation and downstream performance is fairly complex, and merits further investigation. Fig.˜2 of App.˜C shows similar results for cosine similarity instead of squared error.
@@ 4.2 Study 2: Distributed Pretraining
Dist. Training (DiLoCo) - 18.000 189 5818.000\,189\,58 12.767 859 412.767\,859\,4
Dist. ReLoRA (LTE) - 20.973 187 5520.973\,187\,55 13.715 583 3613.715\,583\,36
Identical Random ${\mathbf{P}}_{i}={\mathbf{P}}_{j}$ 21.507 394 5121.507\,394\,51 14.279 870 7914.279\,870\,79
Independent Random $\operatorname{\mathbb{E}}[{\mathbf{P}}_{i}{\mathbf{P}}_{j}^{\top}]=\mathbf{0}$ 20.114 808 0120.114\,808\,01 13.655 087 1713.655\,087\,17
Distributed Random ${\mathbf{P}}_{i}{\mathbf{P}}_{j}^{\top}=\mathbf{0}$ 19.813 587 5219.813\,587\,52 13.514 498 9213.514\,498\,92
@@
Our second experiment targets distributed pretraining of LLMs across poorly-connected and resource-constrained workers, which is important for many applications of interest, from federated training of LLMs to scaling up LLMs across data centers that are not co-located, i.e., where techniques like FSDP are not possible. DiLoCo (Douillard et al., 2024) is a recent and effective approach that has workers train independently for some number of iterations using an inner optimizer, and then uses the average change in parameters from each worker as a “pseudo-gradient” on an outer optimizer that updates a global copy of the parameters (i.e., as in federated learning; McMahan et al., 2017; Reddi et al., 2021). This approach has since been scaled up to train 10B LLMs across distributed workers. [^11] See INTELLECT-1 and (Jaghouar et al., 2024).
However, DiLoCo still assumes that each worker has enough memory to perform a full forward/backward pass on the model, i.e., it does not target memory efficiency. A memory-efficient distributed training approach that is of particular interest in light of the equivalence in §˜3 is LoRA-the-explorer (LTE; Huh et al., 2024), which can be seen as an extension of ReLoRA to the distributed setting. LTE has $K$ independent workers train separate LoRA adapters for a small number of local steps, and then performs a global step by averaging the adapters across workers. The globally-averaged adapter is then merged into the base weights, and optimization continues by resetting and traiing the worker-specific LoRA adapters.
We will now describe how the equivalence in §˜3 can be used to derive an improved version of LTE, which trains only one side of the LoRA adapter in each worker, but initializes the frozen side in a worker-aware manner. Consider a one-sided analogue of LTE, where the weight ${\mathbf{W}}^{(g,l)}_{k}$ for the $k$th worker after $g$ global and $l$ local updates is
@@
By §˜3.2, local steps must correspond to training the worker weights using a gradient transformation,
@@
One approach to initialize/reset the frozen side of worker-specific LoRA adapters (i.e., the gradient projections ${\mathbf{P}}_{k}$) is to sample a projection and broadcast it to all workers. However, the GaLore–LoRA duality suggests a different scheme. §˜3.1 shows that training with linear gradient transformations only optimizes a subspace of the full model, namely $\operatorname{range}({\mathbf{S}}^{\top})$ where ${\mathbf{S}}^{\top}={\mathbf{I}}\otimes{\mathbf{P}}^{\top}$ in the GaLore case. This suggests a different approach to distributed LoRA training, wherein the frozen part of each LoRA adapter, ${\mathbf{S}}_{1},\ldots,{\mathbf{S}}_{K}$, is initialized differently, so that the sum of their ranges allows a larger subspace to be trained. For example, we could sample random semi-orthogonal matrices ${\mathbf{P}}_{1},\ldots,{\mathbf{P}}_{K}$ uniformly at random, assign each worker ${\mathbf{S}}_{i}={\mathbf{I}}\otimes{\mathbf{P}}^{\top}_{i}$, and they will likely each cover different portions of the space. An even stronger strategy would be to demand that $\operatorname{range}({\mathbf{S}}^{\top}_{1}),\ldots,\operatorname{range}({\mathbf{S}}^{\top}_{K})$ must be mutually orthogonal, which can be realized by keeping the ${\mathbf{P}}_{i}$’s as semi-orthogonal, but enforcing that ${\mathbf{P}}_{i}{\mathbf{P}}_{j}^{\top}={\mathbf{0}}$ (e.g., by generating a random $m\times m$ orthogonal matrix and having each worker take a different $d\times m$ submatrix.) Intuitively, this ensures that no worker is duplicating the work of another, since their projections are pairwise orthogonal. We experiment with such identical random, independent random, distributed random initialization schemes.
@@
Dist. Training (DiLoCo) 17.809 353 3317.809\,353\,33 18.000 189 5818.000\,189\,58 18.563 059 7818.563\,059\,78
Dist. ReLoRA (LTE) 23.761 633 0623.761\,633\,06 20.973 187 5520.973\,187\,55 19.538 457 2719.538\,457\,27
Identical Random 23.955 952 7223.955\,952\,72 21.507 394 5121.507\,394\,51 20.316 635 6120.316\,635\,61
Independent Random 20.643 903 4620.643\,903\,46 20.114 808 0120.114\,808\,01 19.971 539 4619.971\,539\,46
Distributed Random 20.317 105 4720.317\,105\,47 19.813 587 5219.813\,587\,52 19.655 980 3219.655\,980\,32
@@
Table 4: Results of the distributed pretraining experiments as we vary the rank of the gradient projections and number of workers. For the DiLoCo baseline, we only vary the number of workers, which means that as we increase the number of workers, the DiLoCo baseline can only benefit since we are training on more data without any downside (i.e., a rank restriction). Note that the for the distributed ReLoRA baseline (which is similar to LTE, Huh et al., 2024), we have double the number of trainable parameters as in the one-sided methods.
@@
Results. The results for the main set of experiments are shown in Tab.˜3. We consider two baselines: (i) DiLoCo (Douillard et al., 2024), which has each worker training independently for $500$ steps before computing a pseudo-gradient that is used to update the global parameters using SGD with Nesterov momentum (Sutskever et al., 2013), and (ii) distributed ReLoRA, which is an analog of ReLoRA but adapted to train like DiLoCo, i.e., one trains the LoRA adapter for $500$ steps and defines the adapter weight as the pseudo-gradient for the Nesterov step; this is very close to LTE. [^12] LTE can be seen as using SGD as the optimizer on the pseudo-gradients, but we found this led to worse results in preliminary experiments. Our distributed GaLore experiments make use of random semi-orthogonal projections since the distributed random initialization for it is easy to compute, and does not add significant communication overhead. [^13] Each worker just needs the seed used to sample the orthogonal matrix, and the indices of the rows it will keep. As in §˜4.1, distributed GaLore leads to degradations at both 200M and 1.3B scales compared to the full distributed training baseline (i.e., DiLoCo). However, our distributed random initialization scheme, where workers are “aware” of each, performs well, thus demonstrating the utility of the gradient transformation–adapter duality from §˜3.
@@
Analysis. We perform a study at the 200M scale over how the number of workers and rank affect performance. Intuitively, larger ranks lead to a larger subspace being trained by each worker (and, in the limit, we should recover something akin to DiLoCo when there is no rank reduction), so we would expect performance to improve as we increase the rank. Indeed, the results for this ablation (shown in Tab.˜4) confirm this intuition, likely because DiLoCo benefits from more workers to get a better estimate of the pseudo-gradient for the outer optimizer step. More surprisingly, we find that this gap is largely bridged by ensuring that different gradient transformations are assigned to each worker, with the distributed initialization once again performing the best. It would be interesting to further study how the effectiveness of the distributed initialization scheme changes as we go to more extreme settings (e.g., hundreds of extremely low-rank workers).